OpenStudy (anonymous):

if EF = 4x-16,FG =3x-5, and EG =27 find the values x,EFand EG

OpenStudy (anonymous):

i dont uderstand this

jimthompson5910 (jim_thompson5910):

by the segment addition postulate, we know that EF + FG = EG

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

using substitution we can then say (4x-6) + (3x-5) = 27

jimthompson5910 (jim_thompson5910):

solve for x, then use this value of x to find EF and FG

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats were i get confused

jimthompson5910 (jim_thompson5910):

(4x-6) + (3x-5) = 27 4x-6 + 3x-5 = 27 7x-11 = 27 x = ??

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

question: is EG really 27? or is there more to it?

OpenStudy (anonymous):

no i need to find the values of x ef and eg

jimthompson5910 (jim_thompson5910):

but there's no need to find EG if that's given to be 27 so that's why I'm thinking there's a typo